# Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulations

### v2

@@ 4.4 Mass and energy outflow rate
The mass and energy outflow rates are crucial properties of galactic winds, as they quantify a galaxy’s ability to redistribute baryons and exhibit distinct features in various feedback models. Although many previous simulations have successfully launched galactic outflows in M82-like galaxies, they often struggle to reproduce the observed mass and energy outflow rates of M82. Even recent studies by Schneider et al. (2020) and Schneider & Mao (2024) moderately underestimate the mass outflow rates of the cool and hot phases compared to observational estimates of the wind in M82 (e.g., Leroy et al. 2015; Martini et al. 2018).
@@
We first explore the mass outflow rate of the cool gas. Leroy et al. (2015) derived the mass density function of both HI and $\mathrm{H}_{2}$, which is defined as the mass per unit length, by integrating the mass within ±$\pm 1.5\,\rm{kpc}$ region along the minor axis. To facilitate comparison with their results, we calculate the mass density function of the cool gas at t=30 Myr in our simulations within a 3.0 kpc-wide slit along the minor axis, as shown in the left panel of Figure 14. Furthermore, we measure the mass outflow rate within a narrower 1 kpc-wide slit and compare it to the results of Martini et al. (2018) in the right panel of Figure 14.
The HI+$\mathrm{H}_{2}$ mass distribution and mass outflow rate at $\rm{z}\gtrsim 1000$ pc in simulations using the variable SNe model are approximately 1.5 dex lower than observational estimates. In contrast, simulations with higher energy injection rates, whether from increased initial metallicity or the fixed SNe energy model, can drive cool gas more efficiently, better matching the observed outflow rates below z = 1500 pc. Notably, a higher initial metallicity of $0.5Z_{\odot}$ yields a cool gas outflow rate comparable to observations throughout the simulation volume.
@@
However, Section 3 indicates that simulations with a fixed SNe energy model or higher initial gas metallicity produce more stars than observed in M82’s recent starburst. To simultaneously replicate both the starburst and outflow characteristics, we need to improve the efficiency of coupling SN energy to the ISM. One potential solution is to incorporate the enhanced feedback effects of clustered supernovae (e.g. Gentry et al. 2017; Fielding et al. 2018). In our subsequent work (Li et al. 2024), we explored the impact of clustered SNe on outflow properties. The inclusion of enhanced feedback from clustered SNe successfully reproduced both the total stellar mass of the starburst and the outflow rates observed in M82.
@@
In addition to the amount of SNe energy injected, the gas supply in the nucleus region of the disk also plays an important role in shaping the properties of outflow. Disabling gas return from star particles drastically reduces the mass of cool gas below 1 kpc and decreases the cool gas outflow rate by a factor of 10 throughout the simulation volume. Meanwhile, a $50\%$ gas return fraction yields a cool gas outflow rate comparable to observations above 1.5 kpc, but significantly lower (by a factor of 10) at 0.5 kpc. The explanations for such features are as follows: a substantial amount of gas is accreted by sink/star particles, depleting the gas supply in the starburst core. Initially, this reduces the resistance to the outflow, but eventually limits the available cool gas to fuel the wind, and hence may increase the outflow rate in the early stage, while lowering the outflow rate in the late stage. In addition, the evolution of outflow rates is not synchronous in different simulations, which can be shown by the results below.
@@
$\eta_{E}=\frac{\dot{E}_{out}}{\dot{E}_{SNe}}$ (22)
where $\dot{M}_{out}$ and $\dot{E}_{out}$ are the total mass outflow rate and energy outflow rate respectively. Here, both $\dot{M}_{out}$ and $\dot{E}_{out}$ are evaluated for all thermal phases across the entire radial extent along the major axis, not just the confined region shown in 14. The results are presented in Figure 15. In simulations with SN feedback, the mass loading factors at z = 0.5 kpc and z = 1.0 kpc are approximately 2-5 and 0.8-5, respectively, at the end of the simulation. The current star formation rate in M82 is uncertain, ranging from 2-10 $\rm{M_{\odot}yr^{-1}}$ (e.g. Kennicutt 1998; Leroy et al. 2015). The total mass outflow rate in M82 is estimated to be $\rm{M_{\odot}yr^{-1}}$ at heights of $\rm{z}\sim 1-2$ kpc, including the cold, cool, warm, and hot phases (Martini et al. 2018; Strickland & Heckman 2009). This implies a mass loading factor of 1-10 at z = 1 kpc, which is broadly consistent with our simulation results.
Figure 15: The left column: mass outflow rate (upper) and mass loading factor (bottom) in various simulations at the height of 500, 1000, and 1500 pc as a function of time. Right: energy outflow rate (upper) and energy loading factor (bottom). Note that due to the fast self-quenching, the mass loading factor calculated at later times may be over-estimated.
@@ 5.1 Comparison with previous simulation
The multiphase outflow driven by starburst in M82 has been studied in a number of simulation works in the past two decades (e.g. Cooper et al. 2008; Melioli et al. 2013; Schneider et al. 2020; Schneider & Mao 2024). In comparison to previous studies, the most important improvement in our work is that the star formation process and related feedback are solved self-consistently. More specifically, stars are formed in dense collapsing gas clumps that emerge from the evolution of the gas disc due to gravitational instability. The formation of stars and consequent feedback are handled by the sink particle module. As a result, the stars formed in our simulation are highly coupled with the dense clouds, allowing us to resolve the early phases in the launch of the outflow in great detail. This includes examining how wind bubbles and SN bubbles expand and break through dense clouds where stars are embedded, as well as how isolated bubbles merge and drive the outflow during the early stages.
Moreover, the launch of cool filaments in our simulations shows some notable features that have not been revealed by previous simulations of the galactic winds. In previous studies (e.g. Cooper et al. 2008), the cold filaments in the outflow were thought to result from the interaction between the hot phase and strings of nearly spherical small dense clouds, which are fragments of the giant clouds in the central disk and were drawn out to the outflow. Due to ram pressure, gas in the cold clouds is stripped and entrained into the hot wind, forming filaments. However, in our simulations, the cold and cool filaments in the wind rarely consist of a dense core. Instead, most of them originated as filaments in the turbulent ISM residing in the nuclear region before being pulled into the outflow. These filaments survived the growth and merge of hot bubbles produced by SNe feedback in the starburst region. The reason why there is barely any small-size cloud remaining in the starburst region in our simulations is as follows: first, the star particles are formed around the density peak of dense gas clumps, and hence the stellar feedback would be more effective in destroying the cores of gas clumps; second, the limited resolution in this work cannot resolve clouds with a size of several parsecs.
@@
Our picture of the survival of cool filaments in the starburst region is consistent with previous high-resolution simulations of the destruction of individual GMCs, with a resolution of about 0.01-0.1 pc. Those simulations show that a considerable fraction of the high-density filamentary structures in the GMCs can survive the feedback from newly born massive stars (see, e.g., Kim et al. 2018; Li et al. 2019; Fukushima & Yajima 2021. We argue that it is probable that both the filamentary structures with length of a few to 100 pc and nearly spherical small cloud/clumps with size a few parsecs, harboring cool and cold gas and residing in the nuclear region before the starburst, can survive the feedback from newly born massive stars, and can then be entrained into the multiphase outflow, and both contribute to the cool filaments in the galactic winds.
@@ 5.2 Limitation of this study
We note that there are several limitations in our current work. First, the recipes and parameters adopted in several processes, such as the star formation, various feedback mechanisms, and gas return, have not been well justified by theory or observations and hence are somewhat arbitrary. The properties of the star clusters and the effect of clustered SN would be different if these recipes and parameters were changed. Second, our model of M82 has not incorporated the tidal interactions between the three members in the M81 group, which would lead to the initial conditions in our simulations deviating from the real case to some extent, especially in the central region. However, we expect that this factor may only change the time scale and peak of the starburst moderately and do not have a significant impact on the launch of outflow. Third, the side length of our simulation box is 2 kpc, which is unable to track the development of wind at a larger radius, and the current results also may have suffered from some boundary effect. Lastly, our current simulation only treats the stars formed in previous starbursts as a gravity potential background, while neglecting the feedback from those stars. With a total mass around $3-5\times 10^{9}M_{\odot}$, the older stars could potentially prevent clustering and suppress star formation to a certain degree through type Ia supernovae and IR radiation.
@@
In the second paper of this series, we carry out simulations in a large volume and with a different recipe to couple the SN feedback energy to the neighbouring ISM, and to return the gas in sink/particles to the ISM. We find that a better agreement with the observations on the total stellar mass and outflow rate can be obtained by moderate revisions on the related modules and the inclusion of the enhanced feedback effect because of clustered SNe. More details can be found in Li et al. 2024. However, we find that further improvement is needed to increase the outflow velocity of cool and cold gas in the wind.
@@ 6 Conclusions
    Our simulations successfully reproduce a starburst in the nuclear region, lasting 20-25 Myr with a peak SFR of $20-45\,\rm{M_{\odot}\,yr^{-1}}$. The total stellar mass formed in the starburst mostly ranges from $5\times 10^{8}\,\rm{M_{\odot}}$ to $6\times 10^{8}\,\rm{M_{\odot}}$, which is however moderately higher than the value in M82 inferred from observation. However, reducing the gas return fraction to less than $50\%$ or adopting a lower initial gas disk mass of $2.5\times 10^{9}\,\rm{M_{\odot}}$, can bring the total stellar mass broadly in line with observations. The observed relation between surface star formation rate and gas surface density in our simulations is several times higher than the classical Kennicutt-Schmidt law but is comparable to that of highly active starburst galaxies and young stellar objects.
@@
    The launch of the multiphase galactic-scale wind occurs in two stages. The first stage, lasting about 10 Myr, involves the formation and breakout of a superbubble foam composed of hot, warm, and dense, cool filamentary gas. Initially, combined radiation, stellar wind, and SN feedback create numerous small bubbles. The feedback ionizes and heats low-density ISM while leaving the high-density filaments relatively unaffected. Subsequently, these small bubbles merge to form a large superbubble filled with multiphase gas, including cool filaments. This superbubble eventually breaks out of the disk. The second stage, spanning 10-15 Myr, involves the development of the kpc-scale multiphase outflow. Hot gas expands outward at the highest velocity, followed by warm gas, and then cool gas in filaments. Interactions between the hot phase and cool filaments stretch the filaments. Most of the cool gas in the wind originates directly from the starburst region rather than cooling from the warm and hot phases
@@
    The outflow velocity of cool gas in our simulations is approximately $30-50\%$ of the observed values in M82. Similarly, the velocities of warm and hot gas are about $20-30\%$ lower than observed. When the galactic wind is fully developed, the mass outflow rate of cool gas in simulations with a total injected SN energy of $\sim 4-6\times 10^{57}\rm{erg}$ is comparable to the observations within the range $2000500<z<$ 2000 pc. However, simulations with lower SN energy input result in significantly lower cool gas outflow rates. The mass loading factor in most simulations ranges from 0.8 to 5, consistent with observational estimates for M82. To better match both the outflow velocities and rates observed in M82, enhanced feedback efficiency due to clustered SNe is necessary, while simultaneously reducing the total stellar mass formed in the simulation.

### v4

@@ 4.4 Mass and energy outflow rate
The mass and energy outflow rates are crucial properties of galactic winds, as they quantify a galaxy’s ability to redistribute baryons and exhibit distinct features in various feedback models. Although many previous simulations have successfully launched galactic outflows in M82-like galaxies, they often struggle to reproduce the observed mass and energy outflow rates of M82. Even recent studies by Schneider et al. (2020) and Schneider and Mao (2024) moderately underestimate the mass outflow rates of the cool and hot phases compared to observational estimates of the wind in M82 (e.g., Leroy et al. 2015; Martini et al. 2018).
@@
We first examine the mass outflow rate of cool ($T<2\times 10^{4}\,\rm{K}$) gas, covering the temperature range of molecular, atomic, and ionized hydrogen. Leroy et al. (2015) derived the mass density function of both HI and $\mathrm{H}{2}$, defined as mass per unit length, by integrating the mass within $\pm 1.5\,\rm{kpc}$ along the minor axis. For comparison, we calculate the mass density function of cool gas at $t=30$ Myr in our simulations within a 3.0 kpc-wide slit along the minor axis, shown in the left panel of Figure 18. Note that HI and $\mathrm{H}{2}$ gas have temperatures below 8000 K; thus, the density of the cool ($T<2\times 10^{4}\,\rm{K}$) phase in M82 should be moderately higher than the values in Leroy et al. (2015), since ionized hydrogen typically has lower density than neutral hydrogen. Furthermore, we measure the mass outflow rate within a narrower 1 kpc-wide slit and compare it with the results of Martini et al. (2018), which consider only HI and should be regarded as a lower limit for the cool gas outflow rate, in the right panel of Figure 18. Because the temperature ranges in our simulations differ from those in Leroy et al. (2015) and Martini et al. (2018), readers should interpret these comparisons with caution.
[…]
@@
On the other hand, Section 3 indicates that simulations with a fixed SNe energy model or higher initial gas metallicity produce a starburst higher than estimation of M82’s recent starburst by $\sim 30\%-100\%$. To simultaneously replicate both the starburst and outflow characteristics, we need to improve the efficiency of coupling SN energy to the ISM. One potential solution is to incorporate the enhanced feedback effects of clustered supernovae (e.g. Gentry et al. 2017; Fielding et al. 2018). In our subsequent work (Li et al. 2025), we explored the impact of clustered SNe on outflow properties. The inclusion of enhanced feedback from clustered SNe successfully reproduced both the total stellar mass of the starburst and the outflow rates observed in M82.
@@
In addition to the amount of SNe energy injected, the gas supply in the nucleus region of the disk also plays an important role in shaping the properties of outflow. Disabling gas return from star particles drastically reduces the mass of cool gas below 1 kpc and decreases the cool gas outflow rate by a factor of 10 throughout the simulation volume. Meanwhile, a $50\%$ gas return fraction yields a cool gas outflow rate comparable to observations above 1.5 kpc, but significantly lower (by a factor of 10) at 0.5 kpc. One potential explanation is that a substantial portion of gas is accreted by sink/star particles, depleting the gas supply in the starburst core. Initially, this reduces resistance to the outflow, but over time it limits the availability of cool gas to fuel the wind and suppresses star formation after $t=10$ Myr (see Figure 3), thereby reducing the total number of SNe injected into the ISM. Consequently, disabling gas return can enhance the outflow rate in the early stage while suppressing it at later times. In addition, the evolution of outflow rates is not synchronized in different simulations, which can be shown by the results below.
@@
$\eta_{E}=\frac{\dot{E}_{out}}{\dot{E}_{\rm{SNe}}}$ (25)
where $\dot{M}_{out}$ and $\dot{E}_{out}$ are the total mass outflow rate and energy outflow rate respectively. Here, both $\dot{M}_{out}$ and $\dot{E}_{out}$ are evaluated for all thermal phases across the entire radial extent along the major axis, not just the confined region shown in Figure 18. The results are presented in Figure 19. In simulations with SN feedback, the mass loading factors at z = 0.5 kpc and z = 1.0 kpc are approximately 2-5 and 0.8-5, respectively, at the end of the simulation. The current star formation rate in M82 is uncertain, ranging from 2-10 $\rm{M_{\odot}yr^{-1}}$ (e.g. Kennicutt 1998; Leroy et al. 2015). The total mass outflow rate in M82 is estimated to be $\sim 20\,\rm{M_{\odot}yr^{-1}}$ at heights of $\rm{z}\sim 1-2$ kpc, including the cold, cool, warm, and hot phases (Strickland and Heckman 2009; Leroy et al. 2015; Martini et al. 2018). This implies a mass loading factor of 1-10 at z = 1 kpc, which is broadly consistent with our simulation results. Note that the nSN simulation has never produced a detectable outflow.
Figure 19: Left: mass outflow rate (top) and mass loading factor (bottom) from various simulations, measured at the height of 500, 1000, and 1500 pc as a function of time. Right: energy outflow rate (top) and energy loading factor (bottom). No detectable galactic wind is produced in nSN, with all corresponding values remaining below the lower limit. Note that, owing to the rapid decline in the star formation rate, the calculated mass loading factor at later times may be overestimated.
@@ 5.1 Comparison with previous simulation
The multiphase outflows driven by starbursts in M82-like galaxies have been extensively investigated over the past two decades (e.g., Cooper et al. 2008; Melioli et al. 2013; Schneider et al. 2020; Schneider and Mao 2024). The principal improvement of this study is the self-consistent treatment of star formation and feedback. In our framework, stars form within dense, gravitationally collapsing clumps that naturally emerge from the evolution of the galactic disc. Their formation and subsequent feedback are modeled with the sink particle method, directly linking star formation to the underlying evolution of gas disc. In contrast, earlier studies generally prescribed the starburst properties, such as intensity, temporal evolution, spatial distribution, and SN feedback, through imposed assumptions rather than outcome of disc evolution.
For example, Cooper et al. (2008) adopted fixed injection rates of $10^{42}$ $\rm erg/s$ and 11 $\rm M_{\odot}/yr$, distributing energy and mass according to the initial gas density to approximate star clusters. Melioli et al. (2013) also used a fixed energy injection rate of $10^{42}$ $\rm erg/s$, but varied the number of SSCs and the duration of star formation. In contrast, Schneider et al. (2020) employed a variable SN injection rate tied to an idealized starburst history: stars formed at 2020 $\rm M_{\odot}/yr$ from 5–35 Myr and 55 $\rm M_{\odot}/yr$ from 35–70 Myr, with clusters of $10^{7}$ $\rm M_{\odot}$ confined within $r<1$ kpc. Their model produced a total feedback energy of $\sim 6\times 10^{57}$ erg, corresponding to an average of $\sim 6\times 10^{42}$ $\rm erg/s$ during 5–35 Myr. More recently, Schneider and Mao (2024) revised the SFH to a distributed burst model, allowing for broader cluster distributions and variable masses, while maintaining a comparable average injection rate of $\sim 6\times 10^{42}$ $\rm erg/s$ over 40 Myr.
By $t=30$ Myr, when outflows are well developed, our simulations form $3.5\times 10^{8}$$6.0\times 10^{8}$ $\rm M_{\odot}$ of stars, with the lower end consistent with observational estimates of $2.0\times 10^{8}$ to $3.5\times 10^{8}$ $\rm M_{\odot}$ (Rieke et al., 1993; Schreiber et al., 2003). The associated SN energy input is $2.5\times 10^{57}$ to $4.3\times 10^{57}$ erg, yielding an average injection rate of $2.6\times 10^{42}$ to $4.5\times 10^{42}$ erg/s. This is a few times higher than the rates in Cooper et al. (2008) and Melioli et al. (2013), but only $40\%$ to $75\%$ of those adopted by Schneider et al. (2020) and Schneider and Mao (2024).
A key advantage of our approach is that stars form naturally within dense gas clouds, rather than being imposed externally. This allows us to resolve the early stages of outflow launching in greater detail, capturing how stellar wind and SN bubbles expand, interact with, and eventually break out of their dense birth environments. Our simulations also reveal how isolated bubbles merge to drive the large-scale flow and how SNe ejecta mix with and enrich the surrounding multiphase ISM. In addition, by incorporating gas return from GMCs, we provide a more self-consistent description of ISM evolution.
Moreover, the development of cool filaments in our simulations shows some notable features that have not been revealed by previous simulations of the galactic winds. In a number of previous studies (e.g. Cooper et al. 2008), cold filaments in the outflow were attributed to interactions between the hot phase and small, dense clouds, which are fragments of giant molecular clouds in the central disk, that were entrained into the wind. These clouds were initially non-filamentary, but subsequent mixing with the hot flow both accelerated them through momentum transfer and stretched them into filamentary structures (Gronke and Oh, 2018; Fielding and Bryan, 2022).
In our simulations, consistent with many previous studies, most of the cool gas in the outflow originates from the disk. However, unlike several earlier works, we find that most of the cool gas in filaments was initially part of filamentary ISM structures in the disk. These pre-existing filaments survive the early stages of supernova feedback, are subsequently entrained into the outflow, and are gradually accelerated through momentum transfer during mixing with the hot wind, all while largely retaining their filamentary morphology.
As shown in Figures 6 and 7, numerous cool filaments and several dense cores within the disk remain less affected by SN feedback than the diffuse ISM, enabling them to survive for 10–15 Myr rather than being rapidly heated or ionized. During this period, SN-driven bubbles in the starburst region expand and merge to form a superbubble, after which filamentary structures are expelled from the disk at $t\sim 15$ Myr, following the outflow of the hot and warm phases. Figures 9 and 10 further illustrate how these filaments escape the disk, interact with other phases in the interface region (z=200z=200–500 pc), and evolve into web-like networks and $\rm H\alpha$-emitting arcs, features reminiscent of those observed by Lopez et al. (2025). Our results highlight survival and transport as the dominant pathway.
@@
Our picture of the survival of cool filaments in the starburst region is consistent with previous high-resolution simulations of the destruction of individual GMCs, with a resolution of about 0.01-0.1 pc. Those simulations show that a considerable fraction of the high-density filamentary structures in the GMCs can survive the feedback from newly born massive stars (see, e.g., Kim et al. 2018; Li et al. 2019; Fukushima and Yajima 2021).
Furthermore, our quantitative analysis (Figure 14) shows the net transfer of material from hotter phases into the cool phase provides only a secondary contribution to the total cool gas mass in the outflow. This net transfer occurs primarily within the interface region between the gas disc and the outflow (200 pc <z< 500 pc) and the uppermost zone. In contrast, the dominant net source of cool gas in the outflow is the transport of pre-existing cool ISM from the central disc region. Nevertheless, mass exchange between phases is still possible: material can condense from hotter phases into the cool phase, but this gain is likely counterbalanced by heating, ionization, or mixing-driven transfer of cool gas into warmer phases.
Nevertheless, we speculate that, in addition to filamentary structures spanning tens to 100 pc, compact, roughly spherical gas clumps, only a few to ten parsecs in size and containing cool or cold gas in the nuclear region prior to the starburst, could also survive stellar feedback and become entrained in the multiphase outflow, as shown in previous simulations. However, our simulations show almost no such small clouds remaining in the starburst region. This likely reflects two factors: (i) star particles form near the density peaks of clumps, where feedback most effectively disrupts their compact cores, and (ii) our resolution is insufficient to fully resolve clouds with only a few parsecs across. Consequently, the contribution of compact, dense clouds to cool filament formation in the wind may be underestimated in our simulations.
@@ 5.2 Limitation of this study
We note that there are several limitations in our current work. First, the recipes and parameters adopted in several processes, such as the star formation, various feedback mechanisms, and gas return, are not yet firmly constrained by theory or observations, and therefore require further verification. The properties of the star clusters and the effect of clustered SN would be different if these recipes and parameters were changed. Second, our model of M82 has not incorporated the tidal interactions between the three members in the M81 group. This omission may cause the initial conditions in our simulations to deviate from reality, particularly in the central region. Third, the initial mass model is derived from present-day observations and assumes a fully developed ISM at the onset of the starburst. Consequently, the adopted initial conditions may not reflect the actual pre-burst environment. We have shown that the initial mass and metallicity moderately influence both the star formation history (SFH) and outflow properties. These factors are expected to alter the timescale and peak of the starburst, as well as the quantitative properties of the outflow, but not the overall evolutionary picture. Fourth, the side length of our simulation box is 4 kpc, which is unable to track the development of wind at a larger radius, and the current results also may have suffered from some boundary effect.
Resolution is another important factor. Higher resolution would allow us to capture faster fragmentation in dense star-forming clumps, resolve pc scale dense cores and the free expansion phase of individual SNRs, and reduce artificial cooling. Our convergence tests show that lower resolution delays the rise of the SFR, slightly prolongs the starburst, and affects the formation of cold and warm structures such as filaments and arcs. As demonstrated by Cooper et al. (2009b), the number of fragments formed from cloud break-up increases with resolution. Thus, at higher resolution, the properties of filamentary and arc-like features in the outflow may differ from our current results to some extent. A higher-resolution study is therefore needed to confirm our conclusions. Lastly, our current simulation only treats the stars formed in previous starbursts as a gravitational potential background, while neglecting the feedback from those stars. With a total mass around $3-5\times 10^{9}M_{\odot}$, the older stars could potentially prevent clustering and suppress star formation to a certain degree through type Ia supernovae and IR radiation.
@@
In the second paper of this series, we carry out simulations in a large volume and with a different recipe to couple the SN feedback energy to the neighbouring ISM, and to return the gas in sink/particles to the ISM. We find that a better agreement with the observations on the total stellar mass and outflow rate can be obtained by moderate revisions on the related modules and the inclusion of the enhanced feedback effect because of clustered SNe. More details can be found in Li et al. (2025). However, we find that further improvement is needed to increase the outflow velocity of cool and cold gas in the wind.
@@ 6 Conclusions
    Our simulations successfully reproduce a starburst in the nuclear region, lasting 20-25 Myr with a peak SFR of $20-45\,\rm{M_{\odot}\,yr^{-1}}$. The total stellar mass formed in the starburst mostly ranges from $5\times 10^{8}\,\rm{M_{\odot}}$ to $6\times 10^{8}\,\rm{M_{\odot}}$, which is however moderately higher than the value in M82 inferred from observation. However, reducing the gas return fraction to less than $50\%$ or adopting a lower initial gas disk mass of $2.5\times 10^{9}\,\rm{M_{\odot}}$, can bring the total stellar mass broadly in line with observations. The observed relation between surface star formation rate and gas surface density in our simulations is several times higher than the classical Kennicutt-Schmidt law but is comparable to that of highly active starburst galaxies.
@@
    The launch of the multiphase galactic-scale wind in our simulations occurs in two stages. The first stage, lasting about 10 Myr, involves the formation and breakout of a superbubble foam composed of hot, warm, and cool filamentary gas. Initially, combined radiation, stellar wind, and SN feedback create numerous small bubbles. The feedback ionizes and heats low-density ISM, while leaving many intermediate to high density filaments relatively intact. As the bubbles expand and merge, they form a large superbubble that contains multiphase gas, including embedded cool filaments, which eventually breaks out of the disk. The second stage, spanning 10-15 Myr, involves the development of the kpc-scale multiphase outflow. Hot gas expands outward at the highest velocity, followed by warm gas, and then cool gas in filaments. Interactions between the hot phase and cool filaments stretch the filaments. Transport of pre-existing cool ISM from the starburst region is the primary net contributor to the total mass of the cool phase in the outflow. In contrast, mass transfer from hotter phases, such as condensation in the boundary layer or precipitation driven by radiative cooling, appears to be largely offset by the concurrent transfer of material from the cool phase to warmer phases during mixing. As a result, these processes make only a minor net contribution to the total cool mass in the outflow, acting primarily within the interface region between disc and outflow (z=200z=200–500 pc) and the uppermost vertical zones ($z>1500$ pc).
@@
    The outflow velocity of cool gas in our simulations is approximately $30-50\%$ of the observed values in M82. Similarly, the velocities of warm and hot gas are about $20-30\%$ lower than observed. When the galactic wind is fully developed, the mass outflow rate of cool gas in simulations with a total injected SN energy of $\sim 4-6\times 10^{57}\rm{erg}$ is comparable to the observations within the range $500<z<2000$ pc. However, simulations with lower SN energy input result in significantly lower cool gas outflow rates. The mass loading factor in most simulations ranges from 0.8 to 5, consistent with observational estimates for M82. To better match both the outflow velocities and rates observed in M82, enhanced feedback efficiency due to clustered SNe is necessary, while simultaneously reducing the total stellar mass formed in the simulation.